Database last updated on Mon, 02 Feb 2026 06:40:53 +0000 by localhost

Instrument constant changes for the week before 2024027 (2 sec, 0 rows)

HG burned out? (AP results when HG is ON) (0 sec, 12 rows)

Hg with step corrections larger than 10 steps or multiple corrections larger than 2.5 steps (0 sec, 1 rows)

SerialYYYYJJJHG largemax correction
085202402720347.9

HP with step corrections larger than 20 steps (0 sec, 1 rows)

SerialYYYYJJJHP largemax correction
0852024027138.0

Non-critical warnings from b-files' comments for days between 2024027 and 2024027 (1 sec, 4 rows)

TypeCommentcountpriorityfnamelocationserialtimevalueminmax
dp:Doing dispersion with internal HG.199B02724.085TORON#08518:50:001-1
ze:Zenith discrepancy = 76499B02724.085TORON#08519:30:4176-1010
dsp:Doing dispersion test with Hg (internal lamp)199B02724.085TORON#08520:32:381-1
checkhg:Last HG had to adjust the micrometer by at least 2199B02724.085TORON#08521:54:131-1

Measurements and tests summary.


_jday_loc_serial_skcdtedhghprssl
027TORONTO#085inside,menu,Saving backup OP_ST.000,Starting schedule inside,Waiting until ZA = 170 now 136.2,Waiting until ZA = 170 now 137.9,Waiting until ZA = 170 now 139.5,Waiting until ZA = 170 now 141.1,Waiting until ZA = 170 now 142.6,Waiting until ZA = 170 now 144.1,Waiting until ZA = 170 now 145.6,Waiting until ZA = 170 now 146.9,Waiting until ZA = 170 now 148.2,Waiting until ZA = 170 now 149.5,Waiting until ZA = 170 now 150.6,Waiting until ZA = 170 now 151.6,Waiting until ZA = 170 now 152.4,Waiting until ZA = 170 now 153.2,Waiting until ZA = 170 now 153.7,Waiting until ZA = 170 now 154.2,Waiting until ZA = 170 now 154.412524111
(1 sec)

Dead time (0 sec)

File nameCountSerialExpected (+/-5) Measured 30-day history
B02724.0851#0852835